OpenStudy (anonymous):

Solve this exponential equation: 2^x-5 = 64^ 3x+1

OpenStudy (anonymous):

confused about the 64^3x+1 is it supposed to be that or 64^3?

hero (hero):

\[ \\2^{x-5} = 64^{3x + 1} \\2^{x-5} = 2^{6^{3x + 1}} \\2^{x-5} = 2^{6(3x + 1)} \\ x - 5 = 6(3x + 1) \\x - 5 = 18x + 6 \\-5 - 6 = 18x - x \\-11 = 17x \\-\frac{11}{17} = x \]